Circular No. 8918

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMETS C/2007 S11, C/2007 U7–U13, C/2007 V3–V10 (SOHO) Further to IAUC 8914, additional near-sun comets have been found

on SOHO website images, as tabulated below; all were Kreutz sungrazers except for C/2007 S11 (no known group affiliation) and the Meyer-group objects C/2007 U7 and C/2007 V10. C/2007 S11 was tiny, stellar in appearance, and very faint (mag 7.5–8). C/2007 U7, C/2007 U11, C/2007 V5, and C/2007 V10 were also tiny, stellar in appearance, and faint (mag ∼ 7.5); C/2007 V8 was described similarly by K. Battams (but with mag ∼ 8). C/2007 V3 was very diffuse and elongated (mag ∼ 7.5). C/2007 V6 was very faint (mag ∼ 8), very diffuse, and elongated. C/2007 V7 was extremely faint (mag ∼ 8.5) and very diffuse. The other seven objects all appeared stellar in C3 images (C/2007 U8 was of mag ∼ 6.5, C/2007 U9 was of mag ∼ 7.5, C/2007 U10 was of mag ∼ 7, C/2007 U12 was of mag ∼ 6, C/2007 U13 and C/2007 V9 were of mag ∼ 6.5, and C/2007 V4 was of mag ∼ 5.5); in C2 images, C/2007 U8 was elongated and somewhat diffuse, C/2007 V4 had a faint thin tail ∼ 23′ long at 5.8 R_{⊙} on Nov. 3.560 UT, and C/2007 V9 was mostly stellar in appearance with a hint of a short faint tail — while the remaining four comets were slightly diffuse (C/2007 U9 being described as faint, and the other three objects showing a hint of a faint tail). New finder code: ZJ = Z. Jin (cf. IAUC 8365).

Comet 2007 UT α_{2000} δ_{2000} Inst. F MPEC
C/2007 S11 Sept. 28.604 12^{h}19.^{m}9 − 0^{o}14′ C2  RK 2008-A71
C/2007 U7 Oct. 27.226 14 02.1 −11 09 C2  HS 2008-A71
C/2007 U8  27.638 13 53.3 −14 36 C3/2 HS 2008-A71
C/2007 U9  28.154 13 59.9 −14 31 C3/2 HS 2008-A71
C/2007 U10  28.679 13 59.6 −14 53 C3/2 HS 2008-A72
C/2007 U11 29.061 14 06.5 −14 28 C2 HS 2008-A72
C/2007 U12  30.179 14 00.6 −15 53 C3/2 TH 2008-A72
C/2007 U13  30.763 14 04.9 −15 59 C3/2 MU 2008-A72
C/2007 V3 Nov. 1.546 14 19.9 −15 38 C2  RM 2008-A75
C/2007 V4  2.471 14 07.1 −17 57 C3/2 AK 2008-A75
C/2007 V5  3.576 14 28.5 −16 20 C2  RK 2008-A75
C/2007 V6  4.643 14 33.4 −16 46 C2  HS 2008-A75
C/2007 V7  5.671 14 37.1 −17 02 C2  RK 2008-A76
C/2007 V8  6.143 14 39.6 −17 16 C2  HS 2008-A76
C/2007 V9  6.196 14 23.8 −19 00 C3/2 RM 2008-A76
C/2007 V10  9.338 14 53.1 −15 31 C2  ZJ 2008-A76
2008 February 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green